OpenStudy (waheguru):

I know that the first energy level has 1 s sub-level and the second energy level has 2s and 2p but then how does the third energy level have an s p and a d because we can only get to 3s and 3p the d's come in the fourth energy level, do the not? and the fs in the sixth not in the second

OpenStudy (amistre64):

im not that up to date on my chemistry lingo, nor the rules that apply.

OpenStudy (waheguru):

do you know about s p d and f?

OpenStudy (amistre64):

nope

OpenStudy (waheguru):

its ok then

OpenStudy (abb0t):

I don't think I fully understand your question. When you're referring to 3\(^{rd}\) energy level you're referring to the s, p and d-orbital subshell, which can hold a maximum of 10 e\(^-\) (total 18). But, the 4s orbital starts to fill before you start filling in the 3d orbital. So before you even reach the d-sublevel you need to occupy the 4s first. Also, as far as f-orbital, I don't think you need to worry too much about those because they are a bit more complex. But they can occupy 14 electrons (total max of 32 e\(^-\)).

OpenStudy (waheguru):

I understand what you are saying so why do we say that that when n = 3 we have 3 sub levels s p and d but in it we really just have 2 3s and 3p ?

OpenStudy (waheguru):

@abb0t

OpenStudy (abb0t):

the n=3 is telling you that the shell contains 9 orbitals, which can hold a max of 18 electrons. 3s, three 3p, and five 3d orbitals and so forth and hence forth...

OpenStudy (waheguru):

Give me a sec ill show you what is confusing me

OpenStudy (waheguru):

The drawing shows what I was thingking I thought that 1s2 - n =1 2s2 2p6 - n = 2 3s2 3p6 - n = 3 4s2 3d10 4p6 - n = 4

OpenStudy (waheguru):

Is this not correct @abb0t

OpenStudy (abb0t):

You're missing the 1s\(^2\) on the n = 2

OpenStudy (abb0t):

1s\(^2\)2s\(^2\)2p\(^6\)

OpenStudy (waheguru):

So n = 1 has just 1s n=2 just has 2s and 2p n=3 has 3s 3d and 3p right @abb0t

OpenStudy (abb0t):

It also has the 1s\(^2\) but you only consider the 2s\(^2\) and 2p\(^n\) because those are the outter most electrons which are capable of reacting. I am curious, is this for general chemistry or inorganic that you're having trouble understanding this in?

OpenStudy (waheguru):

grade 11 chemistry

OpenStudy (waheguru):

Thanks for your help I think I finally understand now : )

OpenStudy (abb0t):

Yes. There is a term that you can look up called \(\sf \color{red}{shielding~effect}\) that might help you understand this a little better. I don't think it is taught in grade-11 chemistry, but definitely in college level chemistry if you proceed further.